A man is 37 years old his child age is 8, how many years ago was the product of the age is 96.
Answer by hamsanash1981@gmail.com(151)

let the years ago be x
then (37-x)(8-x)=96
substitute the value of x = 1,2,3.... until the LHS satisfies RHS
when x =1 LHS = (37-1)(8-1) => 36*7 => 252 IS NOT EQAUL TO RHS
WHEN X =2 LHS =35*6=> 210 IS NOT EQUAL TO RHS
WHEN X= 3 LHS = 34*5=> 170
WHEN X= 4 LHS = 33*4 => 132
WHEN X =5 LHS = 32*3 => 96 IS EQUAL TO RHS.
therefore five years ago their product of ages is equal to 96.